# A comprehensive survey on deep active learning and its applications in medical image analysis

### v1

@@ 5 Active Learning for Medical Image Analysis
Jin et al. (2023b) 2023 EAAI Dermscopy Skin ISIC 2018 Skin Lesion Segmentation Segmentation
X-ray Chest Jaeger et al. (2013) Lung Segmentation Segmentation
@@
Sadafi et al. (2023) 2023 ISBI Histopathology Breast CAMELYON17 Detection of Cancer Metastasesin Lymph Nodes Classification
@@
Qu et al. (2023) 2023 MICCAI Histopathology Colon NCT-CRC-HE-100K Colorectal Cancer Diagnosis Classification
@@
Chen et al. (2023) 2023 MIDL Histopathology Colon PathMNIST Survival Prediction Classification
CT Abdomen OrganAMNIST Classification of Body Organs Classification
Mircoscopy Peripheral Blood BloodMNIST Cell Type Classification Classification
@@
Lou et al. (2023) 2023 TMI Histopathology Seven Organs TCGA-KUMAR Nuclei Segmentation Segmentation
Breast TNBC Nuclei Segmentation Segmentation
Seven Organs MoNuSeg Nuclei Segmentation Segmentation
5.1 Active Learning in Medical Image Classification
@@
Some common clinical tasks, such as disease diagnosis, cancer staging, and prognostic prediction, can be formulated as medical image classification. Most AL works in medical imaging classification directly employ general methods, such as using class-balancing sampling in §3.3.2 to mitigate the long-tail effect of medical imaging datasets. However, specialized design of AL algorithms is required for certain modalities of medical image classification. For example, the classification of chest X-rays often involves the idea of multi-label. Besides, classifying pathological whole-slide images typically needs to be formulated as a multiple instance learning problem. This section will introduce AL works specifically targeted at classification problems in chest X-rays and pathological whole-slide images.
@@ 5.1.1 Chest X-ray and Multi-label Classification
Chest X-ray examinations are crucial for screening and diagnosing lung, cardiovascular, skeletal, and other thoracic diseases. Computer-aided diagnosis in this domain has been extensively researched, including AL works aimed at reducing annotation costs for physicians. Mahapatra et al. (2021) introduced saliency maps to select informative samples for annotation. To aggregate the per-pixel saliency maps into a single scalar, they explored three different approaches, including computing the kurtosis of the saliency map, utilizing multivariate radiomic features, and combining deep features of autoencoders and clustering. Results demonstrated that the aggregation using deep features performs the best. Nguyen et al. (2021) introduced a gist-set to select samples near the decision boundary. Besides, uncertain samples with high entropy were sent for annotation, while the confident samples were assigned as pseudo-labels. Additionally, they adopted momentum updates to enhance the stability of the sample predictions.
However, multiple diseases and abnormalities often coexist simultaneously in diagnosing chest X-rays. Therefore, multi-label classification has been introduced, allowing each sample to be categorized into multiple classes (Baltruschat et al., 2019). Consequently, AL algorithms for chest X-ray classification must adapt to the multi-label setting. Balaram et al. (2022) modified the EDL-based AL to accommodate the multi-label setting. Specifically, they transformed the Dirichlet distribution in EDL into multiple Beta distributions, each corresponding to one class label. They then calculated the entropy of the Beta distributions as the aleatoric uncertainty of the sample. Additionally, they incorporated semi-supervised methods like MeanTeacher (Tarvainen and Valpola, 2017), VAT (Miyato et al., 2018), and NoTeacher (Unnikrishnan et al., 2021) to further reduce annotation costs. Built upon saliency maps, Mahapatra et al. (2022) further introduced GNN to model the inter-relationships between different labels. In this work, each class was treated as a node in a graph, with the relationships between classes represented as edges. They employed various techniques to aggregate information between different classes.
@@ 5.1.2 Pathological Whole-slide Images and Multiple Instance Learning
Compared to modalities like X-ray, CT, and MRI, pathological whole-slide images (WSIs) provide microscopic details at the cellular level, making them critically important for tasks such as cancer staging and prognostic prediction. However, WSIs are very large, with maximum resolutions reaching $000\times 100,$ 000 × 100 , 000 pixels. To handle these large images for deep learning, WSIs are usually divided into many small patches. Fully supervised methods require annotations for each patch, resulting in high annotation costs. AL can effectively improve annotation efficiency. For instance, in classifying breast pathological images, Qi et al. (2019) used entropy as the uncertainty metric. Uncertain patches were sent for annotation, whereas those with low entropy were given pseudo-labels to assist training.
Nevertheless, pathologists might only provide WSI-level annotations in real-world clinical scenarios. Consequently, a prevailing direction in research is to formulate WSI classification as the weakly-supervised multi-instance learning (MIL) (Qu et al., 2022). In this framework, the entire WSI is viewed as a bag, and patches within each WSI are treated as instances within that bag. A well-trained MIL learner can automatically identify relevant patches based on WSI-level labels, thus significantly reducing annotation costs. For example, a trained MIL classifier can automatically spot related patches by annotating whether or not cancer metastasis is present in a WSI. Nonetheless, task-relevant patches are often outnumbered by irrelevant ones, making MIL convergence more challenging. In MIL-based pathological WSI classification, AL filters out irrelevant patches and selects informative patches for annotation. Qu et al. (2023) found that in addition to patches related to the target (e.g., tumors, lymph nodes, and normal cells), WSIs contain many irrelevant patches (e.g., fat, stroma, and debris). Therefore, they adopted the open-set AL (Ning et al., 2022), in which the unlabeled pool contained both target and non-target class samples. They combined feature distributions with prediction uncertainty to select informative and relevant patches of the target class for annotation. Based on attention-based MIL, Sadafi et al. (2023) adopted MC Dropout to estimate both attention and classification uncertainties of each patch, then sent the most uncertain patches in each WSI for expert annotation.
5.2 Active Learning in Medical Image Segmentation

### v3

@@ 5 Active Learning for Medical Image Analysis
Zhao et al. (2021) 2021 JBHI Dermscopy Skin ISIC 2017 image 2,000 (1,600/400) 600 100 Skin Lesion Segmentation Segmentation
X-ray Hand RSNA Bone Age Dataset image 209 (139/20/50) 10 10 Finger Bone Segmentation Segmentation
Ozdemir et al. (2021) 2021 KBS MRI Lower Extremities in-house slice 36 volumes (25/2/9) 64 32 Musculoskeletal Segmentation Segmentation
volume 1 1
Wu et al. (2021) 2021 MedIA CT Lung CC-CCII volume 962 (7:1:2) - 10 COVID-19 Diagnosis Classification
Zhou et al. (2021c) 2021 MedIA Colonoscopy Colon in-house image 4,000 (2,000/2,000) - - Image Quality Assessment Classification
Colonoscopy Colon in-house image 28,250 (16,300/11,950) Polyp Detection Classification
CT Lung in-house PE candidate 6,255 (3,840/2,415) Pulmonary Embolism Detection Classification
Wang and Yin (2021) 2021 MICCAI Mircoscopy (Synthetic) Bacterial Cells VGG Cell image 150 (50/100) - 10% Cell Counting Keypoint Localization
Histopathology Human Bone Marrow MBM 29 (15/14)
Histopathology Human Adipocyte Cells ADI 150 (50/100)
- Various Tissues & Species DCC 176 (100/76)
Xu et al. (2021) 2021 MICCAI Fundus Eye DRIVE region 40 images (20/20) - 10% patches in selected images Retina Vessel Segmentation Segmentation
OCTA Eye ROSE-1 117 images
Table 3: Surveyed Works of Active Learning related to Medical Image Analysis. “-” stands for such information is not available for not provided by the authors or not in the case.
Year Venues Modality ROIs Dataset Sampling Unit Size (Train/Val/Test) or (Train+Val/Test) Initial Pool Size Budget per Round Clinical Task Technical Task
Zhou et al. (2021b) 2021 MICCAI CT Lung MSD scribble, bounding box, extreme point 96 volumes (64/32) - - Tumor Segmentation Segmentation
Colon 190 volumes (126/64)
Kidney KiTS 19 210 volumes (168/42) Kidney & Tumor Segmentation
Chong et al. (2021) 2021 MICCAIW CT Chest (Rahman et al., 2021) image 15,153 (8:1:1) 10% 10% COVID-19 Diagnosis Classification
Nguyen et al. (2021) 2021 MIDL X-ray Chest in-house image 135,309 (131,030/4,279) 6,550 $\leq$6,550 Diagnosis of Airspace Opacity & Lung Lesion Classification
Chest RSNA Pneumonia 26,684 (24,015/2,669) Diagnosis of Pneumonia Classification
Chest CheXpert 228,595 (224,316/4,279) Detection of Pleural Effusion Classification
Mahapatra et al. (2021) 2021 TMI X-ray Chest ChestX-ray8 image 112,120 (7:1:2) 10% 10% Thoracic Disease Diagnosis Classification
Histopathology Colon GlaS patch 165 images Gland Segmentation Segmentation
Nath et al. (2021) 2021 TMI CT Pancreas MSD volume 281 (221/30/30) 20 5 Pancreas & Tumor Segmentation Segmentation
MRI Hippocampus 263 (163/50/50) 10 1 Hippocampus Segmentation Segmentation
Chen et al. (2021) 2021 TPAMI CT Heart in-house volume 168 (126/42) - - Calcification Level Prediction in Aortic Stenosis Classification
Kothawade et al. (2022b) 2022 AAAI X-ray Chest PneumoniaMNIST image 5,856 (4,708/524/624) - - Pneumonia & Normal Classification Classification
Wang et al. (2022b) 2022 AAAI cyro-ET (simulated) - SHREC’19 image 25,000 (24,000/1,000) 10% 5% Subtomogram Classification Classification
Quan et al. (2022) 2022 CVPR X-ray Head Kaggle image 400 (150/250) - - Cephalometric Landmark Detection Keypoint Localization
Hand Payer et al. (2019) 909 (609/300) Hand Landmark Detection Keypoint Localization
Zhang et al. (2022a) 2022 CVPR MRI Spine in-house image 7,295 (7:1:2) 10% 5% Diagnosis of Metastatic Epidural Spinal Cord Compression Classification
Jin et al. (2022c) 2022 Knowledge-based Systems Dermscopy Skin ISIC 2020 image 33,126 (8:1:1) 0 10%, 30% & 50% Skin Lesion Classification Classification
Jin et al. (2022b) 2022 Knowledge-based Systems Dermscopy Skin ISIC 2018 image 2,594 (3:1:1) 0 40%-60% with step of 2.5% Skin Lesion Segmentation Segmentation
X-ray Chest Jaeger et al. (2013) image 707 (3:1:1) 0 30%-50% with step of 2.5% Lung Segmentation Segmentation
Atzeni et al. (2022) 2022 MedIA MRI Brain SATA boundary pixel 35 volumes - - Brain Structure Segmentation Segmentation
Histology Brain in-house 15 stacks Brain Structure Segmentation Segmentation
Dai et al. (2022) 2022 MedIA MRI Brain BraTS 2019 image 335 patients (260/75) 0.5% 1% Brain Tumor Segmentation Segmentation
MALC image 30 patients (20/10) 6% 6% Brain Structure Segmentation
Zhou et al. (2022) 2022 MedIA CT Lung MSD scribble, bounding box, extreme point 96 volumes (64/32) - - Tumor Segmentation Segmentation
Colon 190 volumes (126/64)
Kidney KiTS 19 210 volumes (168/42) Kidney & Tumor Segmentation
Colonoscopy Colon CVC-ClinicDB 29 sequences (23/3/3) Polyp Segmentation
Nath et al. (2022) 2022 MICCAI CT Liver MSD volume 131 (105/26) 0 5% Liver & Tumor Segmentation Segmentation
Hepatic Vessels 303 (242/61) 2% Hepatic Vessels & Tumor Segmentation
Bai et al. (2022) 2022 MICCAI Wireless Capsule Endoscopy Colon CAD-CAP image 1,812 (4:1) 0 10% Polyp Segmentation Segmentation
Balaram et al. (2022) 2022 MICCAI X-ray Chest Chestx-ray8 image 112,120 (7:1:2) 2% 0.5% Thoracic Disease Diagnosis Classification
5% 1%
Wu et al. (2022d) 2022 MICCAI CT Liver LiTS slice 130 (train) - - Liver & Tumor Segmentation Segmentation
CT, MRI CHAOS CT: 20(train) MRI: 120 (60/60)
CT Sliver07 20 (test)
CT MSD 131 (test)
Kothawade et al. (2022c) 2022 MICCAIW X-ray Chest PneumoniaMNIST image 5,856 (4,708/524/624) 100 10 Pneumonia & Normal Classification Classification
Histopathology Colon PathMNIST 107,180 (89,996/10,004/7,180) 50 500 Survival Prediction
Mircoscopy Peripheral Blood BloodMNIST 17,092 (11,959/1,712/3,421) 228 20 Cell Type Classification
Dermscopy Skin ISIC 2018 2,594 1947 20 Skin Lesion Diagnosis
Fundus Eye APTOS-2019 - 673 40 Diabetic Retinopathy Detection
Aklilu and Yeung (2022) 2022 MLHC Laparoscopy Gallbladder CholecSeg8k image 8,080 (4,640/1,600/1,640) 10 10% Segmentation of Laparoscopic Surgical Images Segmentation
Bernhardt et al. (2022) 2022 Nature Communications X-Ray Chest NoisyCXR image 26.6K - - Thoracic Disease Diagnosis Classification
Li et al. (2022) 2022 TMI Histopathology Prostate PANDA image 11,000 10% 10% Gleason Grading of Prostate Cancer Classification
Wu et al. (2022c) 2022 TMI CT Lung CC-CCII slice 750 20% 5% Lung Lesions Segmentation Segmentation
108,676 COVID-19 and Penunomia Diagnosis Classification
Colonoscopy Colon HyperKvasir image 1000 Polyp Segmentation Segmentation
2,717 Colonoscopic Lesion Classification Classification
Mahapatra et al. (2022) 2022 TMI X-ray Chest ChestXpert image 65,240 patients (7:1:2) 10% 10% Thoracic Disease Diagnosis Classification
Table 3: Methodology summarization of surveyed active learning works.“-” stands for such information is not available for not provided by the authors or not in the case.
Year Venues Modality ROIs Dataset Sampling Unit Size (Train/Val/Test) or (Train+Val/Test) Initial Pool Size Budget per Round Clinical Task Technical Task
Khanal et al. (2023) 2023 arXiv MRI Brain BraTS 2018 slice 5,846 (3,673/1,009/1,164) 200 100 Brain Tumor Analysis Classification & Segmentation
X-Ray Chest COVID-QU-Ex image 5,826 (3,728/932/1,166) 100 100 COVID-19 Diagnosis Classification
Chen et al. (2023a) 2023 arXiv Dermscopy Skin Fed-ISIC image 21,989 (17,591/4,398) 500 500 Skin Lesion Diagnosis Classification
@@
Dermscopy Histopathology Fed-Camelyon 455,954 (364,761/91,193) Detection of Cancer Metastasesin Lymph Nodes
Colonoscopy Colon Fed-Polyp 2,187 (1,751/436) 50 50 Polyp Segmentation Segmentation
MRI Prostate Fed-Prostate 1,867 (1,541/326) 20 20 Prostate Segmentation
Fundus Eye Fed-Fundus 1,060 (849/211) Retina Vessel Segmentation
Wang et al. (2023) 2023 arXiv MRI Nose in-house slice 1,057 patients (7:1:2) - 20% Nasopharyngeal Carcinoma Tumor Segmentation Segmentation
Jin et al. (2023b) 2023 EAAI Dermscopy Skin ISIC 2018 image 2,594 patients (3:1:1) 0 10%-30% with step of 2.5% Skin Lesion Segmentation Segmentation
X-ray Chest Jaeger et al. (2013) 704 patients (3:1:1) 0 30%-50% with step of 2.5% Lung Segmentation
Jiménez et al. (2023) 2023 ICCVW Histopathology Gland GlaS patch 165 images (85/16/64) 43% 5% Gland Segmentation Segmentation
Sadafi et al. (2023) 2023 ISBI Histopathology Breast CAMELYON17 WSI 500 0 2 Detection of Cancer Metastasesin Lymph Nodes Classification
Gaillochet et al. (2023b) 2023 MedIA MRI Prostate PROMISE 2012 slice 1,377 (1,020/109/248) 10 10 Prostate Segmentation Segmentation
MRI Hippocampus MSD 9,270 (7,163/350/1,757) Anterior and Posterior Hippocampus Segmentation
Li et al. (2023) 2023 MedIA Ultrasound Breast in-house superpixel 3,200 images (2,600/300/300) 260 130 Breast Tumor Segmenation Segmentation
CT Liver in-house 8,797 images (8,000/400/397) 200 200 Liver Segmenation
Ultrasound Breast BUSI 647 images (400/123/124) 40 40 Breast Tumor Segmenation
X-Ray Chest CXRSet 704 images (500/102/102) 10 10 Lung Segmenation
Bai et al. (2023) 2023 MICCAI CT Liver in-house volumes 941 (752/189) 0 20 Liver Tumor Segmentation Segmentation
Liu et al. (2023a) 2023 MICCAI MRI Heart MSD volume or patch 20 (16/4) 0 3 Left Atrial Segmentation Segmentation
CT Liver 260 (208/52) 5 Liver & Tumor Segmentation
MRI Hippocampus 131 (105/26) Anterior and Posterior Hippocampus Segmentation
CT Pancreas 281 (225/56) Pancreas & Tumor Segmentation
CT Spleen 260 (208/52) Spleen Segmentation
Kadir et al. (2023) 2023 MICCAI OCT Eye Duke superpixel 100 scans (6:2:2) 2% 10% OCT Layer Segmentation Segmentation
AROI 1136 scans (6:2:2) - -
UMN 725 scans (6:2:2) - -
Tang et al. (2023) 2023 MICCAI Ultrasound Carotid CUBS image 3,220 (2,016/1,204) 159 200 Carotid Intima-Media Segmentation Segmentation
in-house 350 (test) - -
@@
Qu et al. (2023b) 2023 MICCAI Histopathology Colon NCT-CRC-HE-100K image 100,000 0 5% Colorectal Cancer Diagnosis Classification
Qiu et al. (2023) 2023 MICCAI Histopathology Breast CAMELYON16 patch 398 WSIs (270/128) - patch size of 4096, 8192, 12288 patch per WSI of 1, 3, 5 Detection of Cancer Metastasesin Lymph Nodes Classification
@@
Chen et al. (2023b) 2023 MIDL Histopathology Colon PathMNIST image 107,180 (89,996/10,004/7,180) 20 10 Survival Prediction Classification
CT Abdomen OrganAMNIST 58,830 (34,561/6,491/17,778) Classification of Body Organs
Mircoscopy Peripheral Blood BloodMNIST 17,092 (11,959/1,712/3,421) Cell Type Classification
Gaillochet et al. (2023a) 2023 MIDL MRI Prostate PROMISE 2012 slice 1,377 (1,020/109/248) 10 10 Prostate Segmentation Segmentation
Qu et al. (2023a) 2023 NeurIPS CT Abdomen AbdomenAtlas-8K volume 8,448 - - Multi-organ Segmentation Segmentation
Lüth et al. (2024) 2023 NeurIPS Dermscopy Skin 25,331 (15,200/3,799/6,332) image 45, 225, 900 45, 225, 900 Skin Lesion Diagnosis Classification
@@
Lou et al. (2023) 2023 TMI Histopathology Seven Organs TCGA-KUMAR patch 30 images (12/4/14) - 5% Nuclei Segmentation Segmentation
Breast TNBC 50 images (30/7/13) 5%
Seven Organs MoNuSeg 44 image (30/14) 7%
Hu et al. (2023) 2023 TMI Histopathology Colon NCT-CRC-HE-100K image 100,000 (4:1) 1%, 2% 1%, 2% Colorectal Cancer Diagnosis Classification
Lung & Colon LC25000 25,000 (4:1) 2% 2% Pulmonary & Colorectal Cancer Diagnosis
Li et al. (2024) 2024 IJCARS MRI Lower Extremities in-house volume or slice 119 volumes (90/9/20) 1 volume or 190 slices 1 volume or 190 slices Musculoskeletal Segmentation Segmentaion
CT 30 volumes (25/1/4) 1 volume or 540 slices 1 volume or 540 slices
Mahapatra et al. (2024) 2024 MedIA X-Ray Chest CheXpert image 224,114 (223,414/200/500) 10 10 Detection of Pleural Effusion Classification
ChestXray14 112,120 (test) Thoracic Disease Diagnosis
Ultrasound Breast BreastMNIST 780 (546/78/156) Breast Cancer Diagnosis
Dermscopy Skin DermaMNIST 10,015 (7,007/1,003/2,005) Skin Lesion Diagnosis
Fundus Eye RetinaMNIST 1,600 (1,080/120/400) Diabetic Retinopathy Severity Grading
Microscopy Kidney TissueMNIST 236,386 (165,466/23,640/47,280) Kidney Cortex Disease Classification
Table 3: Methodology summarization of surveyed active learning works. “-” stands for such information is not available for not provided by the authors or not in the case.
5.1 Active Learning for Medical Image Classification
@@
Common clinical tasks like disease diagnosis, cancer staging, and prognostic prediction can be formulated as medical image classification. Most AL works in medical imaging classification directly employ general methods, such as using class-balancing sampling in §3.3.2 to mitigate the long-tail effect of medical imaging datasets. However, specialized design of AL algorithms is required for certain modalities of medical image classification. For example, the classification of chest X-rays often involves the idea of multi-label. Besides, classifying pathological whole-slide images typically needs to be formulated as a multiple-instance learning problem. This section will introduce AL works specifically targeted at classification problems in chest X-rays and pathological whole-slide images.
@@ 5.1.1 Chest X-ray and Multi-label Classification
Chest X-ray examinations are crucial for screening and diagnosing lung, cardiovascular, skeletal, and other thoracic diseases. Computer-aided diagnosis in this domain has been extensively researched, including AL works aimed at reducing annotation costs for physicians. Mahapatra et al. (2021) introduced saliency maps to select informative samples for annotation. To aggregate the per-pixel saliency maps into a single scalar, they explored three different approaches, including computing the kurtosis of the saliency map, utilizing multivariate radiomic features, and combining deep features of autoencoders and clustering. Results demonstrated that the aggregation using deep features performs the best. Nguyen et al. (2021) introduced a gist-set to select samples near the decision boundary. Besides, uncertain samples with high entropy were sent for annotation, while the confident samples were assigned as pseudo-labels. Additionally, they adopted momentum updates to enhance the stability of the sample predictions. To handle the annotation noise, Bernhardt et al. (2022) proposed a framework called ‘active label cleaning’. This framework ranked samples based on estimated label correctness and labeling difficulty. Experiments on the chest X-ray dataset showed that the proposed method improves performance by efficiently reducing label noise with fewer expert annotations compared to random selection.
However, multiple diseases and abnormalities often coexist simultaneously in diagnosing chest X-rays. Therefore, multi-label classification has been introduced, allowing each sample to be categorized into multiple classes (Baltruschat et al., 2019). Consequently, AL algorithms for chest X-ray classification must adapt to the multi-label setting. Built upon saliency maps, Mahapatra et al. (2022) further introduced GNN to model the inter-relationships between different labels. In this work, each class was treated as a node in a graph, with the relationships between classes represented as edges. They employed various techniques to aggregate information between different classes. As a follow-up work, Mahapatra et al. (2024) further introduced graph multiset transformers (Baek et al., 2020) for more powerful inter-label relationships than GNN.
@@ 5.1.2 Pathological Whole-slide Images and Multiple Instance Learning
Compared to modalities like X-ray, CT, and MRI, pathological whole-slide images (WSIs) provide microscopic details at the cellular level, making them critically important for tasks such as cancer staging and prognostic prediction. However, WSIs are very large, with maximum resolutions reaching $000\times 100,$ 000 × 100 , 000 pixels. To handle these large images for deep learning, WSIs are usually divided into many small patches. Fully supervised methods require patch-level or even cell-level annotations, resulting in high annotation costs. AL can effectively improve annotation efficiency. For instance, in classifying breast pathological images, Qi et al. (2019) used entropy as the uncertainty metric. Uncertain patches were sent for annotation, whereas those with low entropy were given pseudo-labels to assist training. In AL of patch-level histological tissue classification, Hu et al. (2023) proposed category-wise curriculum querying to dynamically adjust the weight of uncertainty sampling of each class. They further proposed negative pre-training with wrong predictions to better distinguish the visually similar classes. To obtain fine-grained cellular annotation from WSI, van der Wal et al. (2021) proposed a human-augmenting AI-based labeling system with the help of AL. An active learner was used to select the next best patch for annotation and a classifier was trained for suggesting annotation. Specifically, Core-Set (Sener and Savarese, 2018) was used as the active learner. Experiments with pathologists demonstrate its ability to reduce workload by around 90% and slightly improve data quality across various cellular labeling tasks.
Nevertheless, pathologists might only provide WSI-level annotations in real-world clinical scenarios. Consequently, a prevailing direction in research is to formulate WSI classification as the weakly-supervised multi-instance learning (MIL) (Qu et al., 2022). In this framework, the entire WSI is viewed as a bag, and patches within each WSI are treated as instances within that bag. A well-trained MIL learner can automatically identify relevant patches based on WSI-level labels, thus significantly reducing annotation costs. For example, a trained MIL classifier can automatically spot related patches by annotating whether or not cancer metastasis is present in a WSI. Nonetheless, task-relevant patches are often outnumbered by irrelevant ones, making MIL convergence more challenging. In MIL-based pathological WSI classification, AL filters out irrelevant patches and selects informative patches for annotation. Based on attention-based MIL, Sadafi et al. (2023) adopted MC Dropout to estimate both attention and classification uncertainties of each patch, then sent the most uncertain patches in each WSI for expert annotation. Qu et al. (2023b) found that in addition to patches related to the target (e.g., tumors, lymph nodes, and normal cells), WSIs contain many irrelevant patches (e.g., fat, stroma, and debris). Therefore, they adopted the open-set AL (Ning et al., 2022), in which the unlabeled pool contained both target and non-target class samples. They combined feature distributions with prediction uncertainty to select informative and relevant patches of the target class for annotation.
5.2 Active Learning for Medical Image Segmentation